# Fast Estimation of Wasserstein Distances via Regression on Sliced Wasserstein Distances

### v1

@@ 3.1 Sliced Wasserstein and Lifted Sliced Wasserstein
where $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ is the projection function, $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is the SWGG (equation 3), and $\sigma\in\mathcal{P}(\mathbb{S}^{d-1})$ is the slicing distribution. Similar to SW, we can obtain variants of PW by choosing $\sigma$. 1. Fixed prior: The original LSW is introduced as in projected Wasserstein (PW) in rowland2019orthogonal which uses the uniform distribution $\mathcal{U}(\mathbb{S}^{d-1})$. 2. Optimization-based: In contrast to the case of one-dimensional projected Wasserstein which is a always a lower bound of Wasserstein distance, SWGG is always an upper bound of Wasserstein distance. Therefore, it is desirable to select $\theta$ that can minimize the corresponding lifted cost, that leads to min SWGG distance: $\text{Min-}SWGG_{p}(\mu,\nu)=\min_{\theta\in\mathbb{S}^{d-1}}\overline{W}_{p}(\mu,\nu;P_{\theta})$. 3. Energy-based: Similar to the case of EBSW, authors in liu2025expected proposes to choose $\sigma$ as an energy-based distribution with the unnormalized density: $p_{\sigma}(\theta)\propto f(-\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})),$ where $f$ is often chosen to be an exponential function with temperature. The authors name the distance as expected sliced transport (EST).
Empirical estimation. For PW, Monte Carlo samples are used to approximate the distance: $\widehat{PW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\frac{1}{L}\sum_{l=1}^{L}\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$ where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$. For Min-SWGG, we can use $\hat{\theta}_{T}$ which is the solution of an optimization algorithm with $T>0$ iterations, e.g., simulated annealing maheyfast2023 , gradient ascent with a surrogate objective maheyfast2023 , and differentiable approximation chapel2025differentiable : $\widehat{\text{Min-}SWGG}_{p}^{p}(\mu,\nu;\hat{\theta}_{T})=\overline{W}_{p}^{p}(\mu,\nu;P_{\hat{\theta}_{T}}).$ For EST, importance sampling estimation is used: $\widehat{EST}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L}))=\sum_{l=1}^{L}\hat{w}_{l}\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$ where $\hat{w}_{l}=\frac{f(-\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}))}{\sum_{l^{\prime}=1}^{L}f(-\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l^{\prime}}})}$ and $\theta_{1},\ldots,\theta_{L}\sim\mathcal{U}(\mathbb{S}^{d-1})$.
Upper bounds. We summarize the connection between PW, Min-SWGG, EST, and Wasserstein distance in the following remark. The connection between Min-SWGG, EST, and Wasserstein distance is discussed in maheyfast2023 ; liu2025expected . The connection between EST and PW can be generalized from the connection between EBSW and SW in nguyen2023energy .
@@ 3.2 Regression of Wasserstein distance on sliced Wasserstein distances
Linear Regression of Wasserstein distance onto SW distances. We now propose linear estimations of Wasserstein distances from SW distances. The motivation is given in Figure 1.
@@
The proof of Proposition 1 in given in Appendix A.1. In practice, we can sample $(\mu_{1},\nu_{1}),\ldots,(\mu_{M},\nu_{M})\sim\mathbb{P}(\mu,\nu)$ to approximate the expectation in equation 10. Let $\hat{\bm{S}}\in\mathbb{R}_{+}^{M\times K}$ be the SW distances matrix i.e., $\hat{\bm{S}}_{ik}=S_{p}^{(k)}(\mu_{i},\nu_{i})$ for $i=1,\ldots,M$, and $\hat{\bm{W}}\in\mathbb{R}_{+}^{M}$ be the Wasserstein distances vector i.e., $\hat{\bm{W}}_{i}=W_{p}(\mu_{i},\nu_{i})$ for $i=1,\ldots,M$, we have the sample-based least-squares estimate: $\hat{\bm{\omega}}_{LSE}=(\hat{\bm{S}}^{\top}\hat{\bm{S}})^{-1}\hat{\bm{S}}^{\top}\hat{\bm{W}}$, which is an unbiased estimate of $\bm{\omega}$.
Remark 3.
From Gauss–Markov theorem johnson2002applied , the least square estimate has the lowest sampling variance within the class of linear unbiased estimators.
Figure 1: Linear regression of the Wasserstein distance on sliced Wasserstein (SW) distances. The left figure illustrates a linear model, interpreted as the $\mathbb{L}_{2}$ projection of the Wasserstein distance onto the linear span of the SW distances. The right figure depicts a special case of a constrained linear model with only two SW distances as predictors, which can be seen as a midpoint method.
@@
$W_{p}(\mu,\nu)=\sum_{k=1}^{K}\omega_{k}SL_{p}^{(k)}(\mu,\nu)+\sum_{k=1}^{K}(1-\omega_{k})SU_{p}^{(k)}(\mu,\nu)+\varepsilon,$ (11)
@@
To estimate $\bm{\omega}=(\omega_{1},\ldots,\omega_{K})$ under the constrained model, we again form the least square estimate, which can be solved using quadratic programming and Monte Carlo estimation. In a special case where $K=1$ i.e., having one lower bound and one upper bound, we can have a closed-form.
Proposition 2.
@@
The proof of Proposition 2 in given in Appendix A.2. The corresponding sample-based estimator for the model is: $\hat{\omega}_{CLSE}=\frac{\frac{1}{M}\sum_{i=1}^{m}(SU_{p}(\mu_{i},\nu_{i})-SL_{p}(\mu_{i},\nu_{i}))(SU_{p}(\mu_{i},\nu_{i})-W_{p}(\mu_{i},\nu_{i}))}{\frac{1}{M}\sum_{i=1}^{M}(SU_{p}(\mu_{i},\nu_{i})-SL_{p}(\mu_{i},\nu_{i})^{2}}.$ Compared to the unconstrained model, the constrained model has half of the parameters. In addition, it adds inductive bias to the model, which is often helpful when having limited observed samples.
Wasserstein Distance Estimation with Few-Shot Regression. We recall that we observe $(\mu_{1},\nu_{1}),\ldots,(\mu_{N},\nu_{N})\sim\mathbb{P}(\mu,\nu)$ in practice. It is not computationally efficient to compute discussed least square estimates using all $N$ pairs of distributions since those estimates require evaluation of Wasserstein distances. We then sample a subset $(\mu_{1}^{\prime},\nu_{1}^{\prime}),\ldots,(\mu_{N}^{\prime},\nu_{M}^{\prime})$ from the original set with $M<<N$. After obtaining an estimate $\hat{\bm{\omega}}$ from $(\mu_{1}^{\prime},\nu_{1}^{\prime}),\ldots,(\mu_{N}^{\prime},\nu_{M}^{\prime})$, we can form estimations of the Wasserstein distances for other pairs and any new pair of distributions given their SW distances.
@@
Extensions on regression. In this work, we focus on regressing the Wasserstein-$p$ distance. If other ground metrics are used e.g., geodesic distances on manifolds, variants of SW distances such as spherical sliced Wasserstein distances bonet2022spherical ; tran2024stereographic ; quellmalz2023sliced , hyperbolic sliced Wasserstein distances bonet2023hyperbolic , sliced Wasserstein for distributions over positive definite matrices bonet2023sliced , and other non-linear variants of sliced Wasserstein bonet2024sliced ; chapel2025differentiable ; tanguy2025sliced ; kolouri2019generalized . However, they might not be upper/lower bounds of the corresponding Wasserstein distances. Moreover, to incorporate uncertainty quantification, we can also perform Bayesian inference box2011bayesian e.g., putting a prior on the regression function.
@@ 4 Experiments
We evaluate our methods in five parts, each with a distinct goal. First, in Section 4.1, we test practical use via $k$-NN on ShapeNetV2, reporting accuracy under different metrics. Second, in Section 4.2, we benchmark RG variants against Wormhole across MNIST point clouds, ShapeNetV2, MERFISH Cell Niches zhang2021spatially , and scRNA-seq atlas persad2023seacells , reporting $R^{2}$/MSE/MAE in low-data regimes. Third, in Section 4.3, we combine our framework with Wormhole to introduce RG-Wormhole, a hybrid that matches Wormhole’s performance while requiring far less training time. We compare training time under varying batch sizes and epochs, as well as embedding, reconstruction, barycenter, and interpolation quality. In Appendix B.1, we run Mixture of Gaussian simulations to verify that our methods approximate the true Wasserstein distance from low to high dimensions. In Appendix B.3, we visualize metric-induced geometry with UMAP mcinnes2018umap . Throughout, $N$ denotes the number of training-set sizes, and $M_{0}$ the number of samples drawn from the training set, yielding $M=\tfrac{M_{0}(M_{0}-1)}{2}$ pairs used to estimate RG coefficients.
@@ 4.1 Point Cloud Classification
We evaluate unconstrained RG variants over a classification task over 10-class ShapeNetV2 with 500 training samples (N=500N{=}500) and estimate RG weights from 10 samples (M=10M{=}10) drawn from the training set. The details of the experimental setting and full results are provided in Appendix B.2.
Results. Table 1 reports $k$-NN accuracy on ShapeNetV2 under different metrics. As expected, WD achieves the best accuracy, with $84.2\%$ at k=5k{=}5. Among single sliced-based metrics, SW and EBSW, are the strongest, though they cap at about $72.5\%$ top-1. Our RG methods close much of the gap to Wasserstein. Both RG-s and RG-e consistently achieve around $82.5\%$ top-1 accuracy with high correlation to Wasserstein ($R^{2}\approx 0.93$). The multi-metric extensions further improve stability: RG-se and RG-seo reach up to $83.5\%$ accuracy with $R^{2}$ as high as $0.96$, essentially matching Wasserstein.
@@
RG-se (constr.) 0.86 $8.4\times 10^{-2}$ $2.2\times 10^{-1}$ 0.92 $1.4\times 10^{-2}$ $9.3\times 10^{-2}$ 0.92 $1.3\times 10^{-5}$ $3.0\times 10^{-3}$ 1.00 $1.3\times 10^{-5}$ $1.0\times 10^{-3}$
RG-seo (constr.) 0.86 $7.8\times 10^{-2}$ $2.2\times 10^{-1}$ 0.92 $1.3\times 10^{-2}$ $9.1\times 10^{-2}$ 0.92 $1.3\times 10^{-5}$ $3.0\times 10^{-3}$ 0.99 $8.2\times 10^{-5}$ $9.0\times 10^{-3}$
@@
We compare our RG framework with Wormhole within the same training sizes, matching the preprocessing of haviv2024wasserstein across four datasets spanning dimensionality: MNIST pixel point clouds (2D), ShapeNetV2 point clouds (3D), MERFISH Niche Cells (254D), and scRNA-seq (2,5002{,}500D). We train on $N\!\in\!\{10,50,100,200\}$ random pairs and evaluate $R^{2}$/MSE/MAE against exact WD. For fairness, the number of training pairs for Wormhole equals the number used to estimate the linear coefficients for RG variants, i.e., M=NM{=}N. Full results appear in Figures 613 with settings in Appendix B.4; Table 2 summarizes the M=100M{=}100 case, and other $M$ follow the same pattern.
@@ 4.3 RG-Wormhole: Accelerating Wormhole with Regression of Wasserstein
We run five experiments of both models to empirically show that RG-Wormhole is much faster than Wormhole while keeping similar effectiveness. First, we measure training time by training Wormhole and RG-Wormhole under the same optimizer and schedule, sweeping batch sizes 4–20 and reporting wall-clock time for training-set sizes $N\!\in\!\{10,50,100,200\}$. Second, we assess encoders via $R^{2}$/MSE/MAE between learned pairwise distances and exact Wasserstein. Third, we evaluate decoders via the Wasserstein loss between each input shape and its reconstruction. Fourth, we examine barycenters by decoding each class’s mean embedding and visualizing results. Finally, we study interpolation by decoding linear paths between two embeddings and visualizing the trajectories. Across all experiments, hyperparameters match Wormhole; the only change in RG-Wormhole is replacing every use of the Wasserstein distance in the encoder and decoder losses with the calibrated unconstrained RG variants. For RG-Wormhole, we estimate the RG coefficient using 10 random training samples (M=10M{=}10) before plugging into Wormhole. We provide some results in Figures23 though the details of experimental settings and full results can be found in Appendix B.5.

### v2

@@ 3.1 Sliced Wasserstein and Lifted Sliced Wasserstein
where $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ is the projection function, $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is the SWGG (equation 3), and $\sigma\in\mathcal{P}(\mathbb{S}^{d-1})$ is the slicing distribution. Similar to SW, we can obtain variants of PW by choosing $\sigma$.
1. Fixed prior: The original LSW is introduced as in projected Wasserstein (PW) in Rowland et al. (2019), which uses the uniform distribution $\mathcal{U}(\mathbb{S}^{d-1})$.
2. Optimization-based: In contrast to the case of one-dimensional projected Wasserstein, which is always a lower bound of Wasserstein distance, SWGG is always an upper bound of Wasserstein distance. Therefore, it is desirable to select $\theta$ that can minimize the corresponding lifted cost, that leads to min SWGG distance:
$\text{Min-}SWGG_{p}(\mu,\nu)=\min_{\theta\in\mathbb{S}^{d-1}}\overline{W}_{p}(\mu,\nu;P_{\theta})$
3. Energy-based: Similar to the case of EBSW, authors in Liu et al. (2025) proposes to choose $\sigma$ as an energy-based distribution with the unnormalized density: $p_{\sigma}(\theta)\propto f(-\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})),$ where $f$ is often chosen to be an exponential function with temperature. The authors name the distance as expected sliced transport (EST).
Empirical estimation. For PW, Monte Carlo samples are used to approximate the distance: $\widehat{PW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\frac{1}{L}\sum_{l=1}^{L}\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$ where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$. For Min-SWGG, we can use $\hat{\theta}_{T}$ which is the solution of an optimization algorithm with $T>0$ iterations, e.g., simulated annealing (Mahey et al., 2023), gradient ascent with a surrogate objective (Mahey et al., 2023), and differentiable approximation (Chapel et al., 2025): $\widehat{\text{Min-}SWGG}_{p}^{p}(\mu,\nu;\hat{\theta}_{T})=\overline{W}_{p}^{p}(\mu,\nu;P_{\hat{\theta}_{T}}).$ For EST, importance sampling estimation is used:
$\widehat{EST}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L}))=\sum_{l=1}^{L}\hat{w}_{l}\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$
where $\hat{w}_{l}=\frac{f(-\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}))}{\sum_{l^{\prime}=1}^{L}f(-\overline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l^{\prime}}})}$ and $\theta_{1},\ldots,\theta_{L}\sim\mathcal{U}(\mathbb{S}^{d-1})$.
Upper bounds. We summarize the connection between PW, Min-SWGG, EST, and Wasserstein distance in the following remark. The connection between Min-SWGG, EST, and Wasserstein distance is discussed in Mahey et al. (2023); Liu et al. (2025). The connection between EST and PW can be generalized from the connection between EBSW and SW in Nguyen & Ho (2023).
@@ 3.2 Regression of Wasserstein distance on sliced Wasserstein distances
Linear Regression of Wasserstein distance onto SW distances. We now propose linear estimations of Wasserstein distances from SW distances.
@@
The detail of Remark 3 in given in Appendix A.1. In practice, we can sample $(\mu_{1},\nu_{1}),\ldots,(\mu_{M},\nu_{M})\sim\mathbb{P}(\mu,\nu)$ to approximate the expectation in equation 10. Let $\hat{\bm{S}}\in\mathbb{R}_{+}^{M\times K}$ be the SW distances matrix i.e., $\hat{\bm{S}}_{ik}=S_{p}^{(k)}(\mu_{i},\nu_{i})$ for $i=1,\ldots,M$, and $\hat{\bm{W}}\in\mathbb{R}_{+}^{M}$ be the Wasserstein distances vector i.e., $\hat{\bm{W}}_{i}=W_{p}(\mu_{i},\nu_{i})$ for $i=1,\ldots,M$, we have the sample-based least-squares estimate: $\hat{\bm{\omega}}_{LSE}=(\hat{\bm{S}}^{\top}\hat{\bm{S}})^{-1}\hat{\bm{S}}^{\top}\hat{\bm{W}}$, which is an unbiased estimate of $\bm{\omega}$. It is well-known that the linear model can be seen as $\mathbb{L}_{2}$ projection of the Wasserstein distances vector $\hat{\bm{W}}$ onto the linear span of the SW distances vectors $\hat{\bm{S}}^{(1)},\ldots,\hat{\bm{S}}^{(K)}$. We illustrate the idea in the left figure in Figure 1.
Figure 1: Linear regression of the Wasserstein distance vector $\hat{\bm{W}}$ on sliced Wasserstein (SW) distances $\hat{\bm{S}}^{(1)},\ldots,\hat{\bm{S}}^{(K)}$. The left figure illustrates a linear model, interpreted as the $\mathbb{L}_{2}$ projection of the Wasserstein distance onto the linear span of the SW distances. The right figure depicts a special case of a constrained linear model with only two SW distances as predictors, which can be seen as a midpoint method.
@@
$W_{p}(\mu,\nu)=\frac{1}{K}\sum_{k=1}^{K}\omega_{k}SL_{p}^{(k)}(\mu,\nu)+\frac{1}{K}\sum_{k=1}^{K}(1-\omega_{k})SU_{p}^{(k)}(\mu,\nu)+\varepsilon,$ (11)
@@
To estimate $\bm{\omega}=(\omega_{1},\ldots,\omega_{K})$ under the constrained model, we again form the least square estimate, which can be solved using quadratic programming and Monte Carlo estimation. In a special case where $K=1$, i.e., having one lower bound and one upper bound, we can have a closed-form.
Remark 4.
@@
The detail of Remark 4 in given in Appendix A.2. The corresponding sample-based estimator for the model is: $\hat{\omega}_{CLSE}=\frac{\frac{1}{M}\sum_{i=1}^{m}(SU_{p}(\mu_{i},\nu_{i})-SL_{p}(\mu_{i},\nu_{i}))(SU_{p}(\mu_{i},\nu_{i})-W_{p}(\mu_{i},\nu_{i}))}{\frac{1}{M}\sum_{i=1}^{M}(SU_{p}(\mu_{i},\nu_{i})-SL_{p}(\mu_{i},\nu_{i})^{2}}.$ We show the idea in the right figure in Figure 1. Compared to the unconstrained model, the constrained model has half of the parameters. In addition, it adds inductive bias to the model, which is often helpful when having limited observed samples.
Wasserstein Distance Estimation with Few-Shot Regression. We recall that we observe $(\mu_{1},\nu_{1}),\ldots,(\mu_{N},\nu_{N})\sim\mathbb{P}(\mu,\nu)$ in practice. It is not computationally efficient to compute the discussed least square estimates using all $N$ pairs of distributions since those estimates require evaluation of Wasserstein distances. We then sample a subset $(\mu_{1}^{\prime},\nu_{1}^{\prime}),\ldots,(\mu_{N}^{\prime},\nu_{M}^{\prime})$ from the original set with $M<<N$. After obtaining an estimate $\hat{\bm{\omega}}$ from $(\mu_{1}^{\prime},\nu_{1}^{\prime}),\ldots,(\mu_{N}^{\prime},\nu_{M}^{\prime})$, we can form estimations of the Wasserstein distances for other pairs and any new pair of distributions given their SW distances.
@@
Extensions on regression. In this work, we focus on regressing the Wasserstein-$p$ distance. If other ground metrics are used e.g., geodesic distances on manifolds, variants of SW distances such as spherical sliced Wasserstein distances (Bonet et al., 2023a; Tran et al., 2024; Quellmalz et al., 2023), hyperbolic sliced Wasserstein distances (Bonet et al., 2023b), sliced Wasserstein for distributions over positive definite matrices (Bonet et al., 2023c), and other non-linear variants of sliced Wasserstein (Bonet et al., 2025; Chapel et al., 2025; Tanguy et al., 2025; Kolouri et al., 2019). However, they might not be upper/lower bounds of the corresponding Wasserstein distances. Moreover, to incorporate uncertainty quantification, we can also perform Bayesian inference (Box & Tiao, 2011), e.g., putting a prior on the regression function.
@@ 4 Experiments
We evaluate our methods in several parts, each with a distinct goal. First, in Section 4.1, we test practical use via $k$-NN on ShapeNetV2, reporting accuracy under different metrics. Second, in Section 4.2, we benchmark RG variants against Wormhole across MNIST point clouds, ShapeNetV2, MERFISH Cell Niches Zhang et al. (2021), and scRNA-seq atlas Persad et al. (2023), reporting $R^{2}$/MSE/MAE in low-data regimes. Third, in Section 4.3, we combine our framework with Wormhole to introduce RG-Wormhole, a hybrid that matches Wormhole’s performance while requiring far less training time. We compare training time under varying batch sizes and epochs, as well as embedding, reconstruction, barycenter, and interpolation quality. In Appendix B.1, we run MoG simulations to verify that our methods approximate the true Wasserstein from low to high dimensions. In Appendix B.3, we visualize metric-induced geometry with UMAP McInnes et al. (2018). In Appendix B.6, we compare RG with classical methods in the many-pairs setting. To illustrate robustness, we further evaluate the method under inter-class and intra-class settings in Appendix B.7. Finally, we investigate whether there are consistent patterns in the optimal RG weights across datasets; see Appendix B.8. Throughout, $N$ denotes the number of training-set sizes, and $M_{0}$ the number of samples drawn from the training set, yielding $M=\tfrac{M_{0}(M_{0}-1)}{2}$ pairs used to estimate RG coefficients.
@@ 4.1 Point Cloud Classification
We evaluate unconstrained RG variants over a classification task over 10-class ShapeNetV2 with 500 training samples (N=500N{=}500) and estimate RG weights from 10 samples ($M_{0}{=}10$) drawn from the training set. The details of the experimental setting and full results are provided in Appendix B.2.
Results. Table 1 reports $k$-NN accuracy on ShapeNetV2 under different metrics. As expected, WD achieves the best accuracy, with $84.2\%$ at k=5k{=}5. Among single sliced-based metrics, SW and EBSW, are the strongest, though they cap at about $72.5\%$ top-1. Our RG methods close much of the gap to Wasserstein. Both RG-s and RG-e consistently achieve around $82.5\%$ top-1 accuracy with high correlation to Wasserstein ($R^{2}\approx 0.9$). The multi-metric extensions further improve stability: RG-se and RG-seo reach up to $83.5\%$ accuracy with $R^{2}$ as high as $0.93$, essentially matching Wasserstein.
@@
RG-se (constr.) 0.84 $9.8\times 10^{-2}$ $2.4\times 10^{-1}$ 0.92 $1.4\times 10^{-2}$ $9.3\times 10^{-2}$ 0.91 $1.5\times 10^{-5}$ $3.0\times 10^{-3}$ 1.00 $2.4\times 10^{-5}$ $2.0\times 10^{-3}$
RG-seo (constr.) 0.85 $9.0\times 10^{-2}$ $2.3\times 10^{-1}$ 0.91 $1.7\times 10^{-2}$ $1.0\times 10^{-1}$ 0.92 $1.3\times 10^{-5}$ $3.0\times 10^{-3}$ 1.00 $2.2\times 10^{-5}$ $2.0\times 10^{-3}$
@@
We compare our RG framework with Wormhole within the same training sizes, matching the preprocessing of (Haviv et al., 2024) across four datasets spanning dimensionality: MNIST pixel point clouds (2D), ShapeNetV2 point clouds (3D), MERFISH Niche Cells (254D), and scRNA-seq (2,5002{,}500D). We train on $N\!\in\!\{10,50,100,200\}$ random pairs and evaluate $R^{2}$/MSE/MAE against exact WD. For fairness, the number of training pairs for Wormhole equals the number used to estimate the linear coefficients for RG variants, i.e., $M_{0}{=}N$. Full results appear in Figures 613 with settings in Appendix B.4; Table 2 summarizes the $M_{0}{=}100$ case, and other $M_{0}$ follow the same pattern.
@@ 4.3 RG-Wormhole: Accelerating Wormhole with Regression of Wasserstein
We run five experiments of both models to empirically show that RG-Wormhole is much faster than Wormhole while keeping similar effectiveness. First, we measure training time by training Wormhole and RG-Wormhole under the same optimizer and schedule, sweeping batch sizes 4–20 and reporting wall-clock time for training-set sizes $N\!\in\!\{10,50,100,200\}$. Second, we assess encoders via $R^{2}$/MSE/MAE between learned pairwise distances and exact Wasserstein. Third, we evaluate decoders via the Wasserstein loss between each input shape and its reconstruction. Fourth, we examine barycenters by decoding each class’s mean embedding and visualizing results. Finally, we study interpolation by decoding linear paths between two embeddings and visualizing the trajectories. Across all experiments, hyperparameters match Wormhole; the only change in RG-Wormhole is replacing every use of the Wasserstein distance in the encoder and decoder losses with the calibrated unconstrained RG variants. For RG-Wormhole, we estimate the RG coefficient using 10 random training samples ($M_{0}{=}10$) before plugging into Wormhole. We provide some results in Figures23 though the details of experimental settings and full results can be found in Appendix B.5.